OpenStudy (anonymous):

Given 28.2 grams of an unknown substance, if the substance absorbs 2165 joules of energy and the temperature increases by 35 Kelvin, what is the specific heat of the substance? Select one: a. 2.14 x 106 J/g·K b. 2.19 x 100 J/g·K c. 5.73 x 10-4 J/g·K d. 2.69 x 103 J/g·K

OpenStudy (anonymous):

what is heat defined as?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

is it kelvin?

OpenStudy (anonymous):

you don't know the equation for specific heat?

OpenStudy (anonymous):

Q = m*c*∆T

OpenStudy (anonymous):

solve that for c, lol.

OpenStudy (anonymous):

idk how to the whole thing confuses me

OpenStudy (anonymous):

use algebra to solve for c in the equation Q=m*c*t

OpenStudy (anonymous):

would it be b

OpenStudy (anonymous):

no wait its d

OpenStudy (anonymous):

help

OpenStudy (somy):

use the formula q= mcT q is energy (in joules) m is mass (in grams) c is specific heat capacity T is temperature (in celsius) ( to convert from Kelvin to Celsius you need to minus 273 from your Kelvin reading) You want to find specific heat capacity so make C as a subject, you'll get c= q/(mT) now just substitute values given and you'll find the answer :)

OpenStudy (somy):

Oh wait your answers are given with Kelvin, so don't convert anything, simply substitute values given

OpenStudy (somy):

I guess i understand what you are confused about. Its those *106 or *100 right? sadly i don't really know it myself, maybe they don't mean anything but are just there to confuse us? I mean i've seen such thing happening a lot but when i'll check out with couple of people and tell you)